# State-space models are accurate and efficient neural operators for dynamical systems

### v1

@@ 2.1 Neural Operators
Neural operators are a special class of machine learning models designed to learn (nonlinear) mappings between infinite-dimensional spaces, making them particularly suitable for efficiently solving partial differential equations (PDEs). The dynamical systems problem under consideration in this paper also falls into this field. Unlike traditional numerical methods that require mesh generation and can be computationally intensive, neural operators learn a representation of the PDE solution operator directly, allowing for fast evaluations on new inputs once trained. This approach provides significant advantages in terms of generalization and scalability, enabling more rapid and flexible solutions across a variety of complex scenarios in physics and engineering.
@@ 3.2 State Space Model (SSM) and Mamba
where $\boldsymbol{\bar{A}}:=\exp(\Delta\cdot\boldsymbol{A})$, $\boldsymbol{\bar{B}}:=(\Delta\cdot\boldsymbol{A})^{-1}(\exp(\Delta\cdot\boldsymbol{A})-I)\cdot\Delta\boldsymbol{B}$, and $\Delta$ is the discretization step size. The recurrence formulas (2) inherently enable the modeling of unbounded context and linear-time inference but they are not conducive to efficient parallel training. Fortunately, equation (2) can be further rewritten using a convolution formula as
@@ 3.3 Transformers
To address the efficiency limitation, several attention variants with a linear complexity havde been recently proposed and studied. In the context of operator learning, a Galerkin-type attention was first introduced by [5] as a linear variant of attention and later employed by [27, 19],
@@ 4 Computational Experiments
We describe the implementation in detail. We generate 10K train and 10K test data with a sensor/discretization size of 100, i.e., the sensors are on $\{0.01,0.02,\cdots,0.99,1\}$, where the input function $u(t)$ and the target function $s(t)/s_{t}(t)$ are discretized. Following the original setting in DeepONet, the input functions $u(t)$ are randomly drawn from a mean zero Gaussian random field (GRF) with a length scale of 0.2, i.e., the input functions follow the distribution $\mathcal{GP}\left(0,k_{l}(x,y)\right)$ where $k_{l}(x,y)=\exp\left(-\frac{(x-y)^{2}}{2l^{2}}\right)$ is the RBF kernel with length scale $l$ and $\mathcal{GP}$ denotes a Gaussian process. The output function $s(t)$ or $s_{1}(t)$ can correspondingly be solved via traditional numerical methods. The train loss and the test metric are all the mean square errors (MSE). We train all models for 10,001 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [26] with a 1e-3 initial learning rate, which decays linearly to zero at the end of the training. We keep the parameter numbers of different models similar for a fair comparison. We run the code for five independent random seeds and report the average.
The model structure hyperparameter settings are summarized below. LSTM, GRU, Mamba, and Mamba2 are one-layer models with 32 hidden dimensions. DeepONet’s trunk and branch nets are all five-layer networks with 24 hidden dimensions. FNO is a three-layer model with 16 hidden dimensions and 8 modes. We choose a one-layer LNO with 16 hidden dimensions and 8 modes. We opt for a four-layer Transformer with 12 hidden dimensions, 48 feedforward dimensions, and 4 heads. For all three Oformers with vanilla, Galerkin, or Fourier attention, we opt for three-layer Oformers with 12 hidden dimensions, 12 feedforward dimensions, and 4 heads. GNOT has three layers, 8 hidden dimensions and 8 feedforward dimensions, 4 heads, and 2 experts.
@@
LSTM 8545 1643MB 31min 1.625E-7 9.509E-6 1.322E-8
GRU 6433 1629MB 30min 3.627E-8 3.215E-6 7.117E-9
DeepONet 7272 1381MB 34min 1.296E-7 2.147E-5 3.500E-8
FNO 7009 1831MB 59min 6.032E-4 1.153E-3 1.245E-4
LNO 6721 1797MB 52min 2.214E-7 4.775E-5 1.073E-8
Transformer 7597 1599MB 119min 3.821E-5 1.349E-4 2.381E-6
Oformer (Vanilla) 8533 1491MB 85min 4.141E-8 1.370E-6 7.800E-9
Oformer (Galerkin) 8461 1391MB 90min 5.439E-8 1.938E-6 4.641E-9
Oformer (Fourier) 8461 1391MB 90min 4.434E-8 1.618E-6 4.527E-9
GNOT 8671 1419MB 220min 1.685E-7 3.845E-6 1.373E-8
Mamba (Ours) 6593 1425MB 35min 3.333E-9 4.351E-8 1.301E-9
Mamba2 (Ours) 6009 1282MB 109min 2.106E-8 2.015E-7 4.464E-8
Table 2: Test MSE error results for 1D dynamical systems, including the anti-derivative operator, nonlinear ODE, and gravity pendulum, from DeepONet’s benchmark cases [32] corresponding to Section 4.1 in this paper. The errors are MSE. The best model is bold, the second best is red, and the third best is blue.
Table 2 shows the computational results, where we report the model name, the number of total parameters in the model, the memory cost in Megabyte (MB) obtained from the command “nvidia-smi” on the Linux GPU server where only the training part is considered into the memory cost evaluation, the total training time in minutes (min), and the model MSE prediction error on the three operator learning tasks. The MSE error metric was initially adopted by Lu et al. [32], and we strictly follow them for a clear comparison.
@@
  • Regarding the proposed Mamba’s performances, Mamba is the best model for all three operator learning tasks. It outperforms the second-best model by a large margin. Mamba2 is generally good and ranks the second best in the anti-derivative and nonlinear ODE operators.
@@
  • We construct models with similar number of parameters. Hence, their memory costs are similar. However, the total training time differs drastically due to different models’ computational complexity. Specifically, the expensive self-attention mechanism makes Transformer-related models (Transformer, Oformer, and GNOT) the slowest, while RNNs and Mamba are relatively faster.
@@ 4.2 Finite Regularity and Discontinuous Solutions
The training loss is MSE, and the test metric is the relative $L_{2}$ error. The test metric is similar with the one used by Shih et al. [45] for a direct comparison. We train all models for 50,001 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [26] with a 1e-3 initial learning rate, which decays linearly to zero at the end of training. We keep the number of parameters of different models similar for a fair comparison. We run the code for five independent random seeds and report the average.
@@
The model structure hyperparameter settings are summarized below. LSTM and GRU are one-layer models with 32 hidden dimensions. DeepONet’s trunk and branch nets are all five-layer networks with 24 hidden dimensions. FNO is a three-layer model with 16 hidden dimensions and 8 modes. We choose a one-layer LNO with 16 hidden dimensions and 8 modes. We opt for a four-layer Transformer with 12 hidden dimensions, 48 feedforward dimensions, and 4 heads. For all three Oformers with vanilla, Galerkin, or Fourier attention, we opt for three-layer Oformers with 12 hidden dimensions, 12 feedforward dimensions, and 4 heads. GNOT has three layers, 8 hidden dimensions and 8 feedforward dimensions, 4 heads, and 2 experts. Regarding Mamba, we opt for a block with skip connection and layer normalization, with 8 model dimensions and 8 intermediate dimensions.
The Izhikevich Model
Model Params Memory Time Rel. $L_{2}$ Err Model Params Memory Time Rel. $L_{2}$ Err
LSTM 8545 1863MB 52min 1.266E-1 Transformer 7597 5121MB 691min 1.448E-1
GRU 6433 1821MB 52min 7.980E-2 Oformer-V 8533 3603MB 491min 2.245E-1
DeepONet 14496 1383MB 45min 2.363E-1 Oformer-G 8461 1611MB 378min 1.531E-1
FNO 7009 1859MB 78min 1.541E-1 Oformer-F 8461 1611MB 378min 1.716E-1
LNO 6721 1973MB 72min 2.936E-1 GNOT 8671 1725MB 791min 1.546E-1
Mamba 9649 1599MB 52min 1.327E-1
Tempered Fractional LIF Model
Model Params Memory Time Rel. $L_{2}$ Err Model Params Memory Time Rel. $L_{2}$ Err
LSTM 8577 1707MB 34min 3.673E-2 Transformer 7609 2357MB 174min 2.695E-2
GRU 6465 1703MB 34min 3.368E-2 Oformer-V 8545 1897MB 121min 1.382E-2
DeepONet 14472 1383MB 34min 2.800E-1 Oformer-G 8473 1457MB 155min 6.079E-2
FNO 7025 1819MB 56min 2.142E-2 Oformer-F 8473 1457MB 155min 4.758E-2
LNO 6737 1859MB 50min 9.113E-1 GNOT 8679 1531MB 335min 5.788E-2
Mamba 9665 1493MB 43min 1.905E-2
@@
Table 3: Results for finite regularity and discontinuous solutions adopted from Shih et al. [45] corresponding to Section 4.2 in this paper. Visualization of the solutions is presented in Figure 2. Here, we present the test relative $L_{2}$ errors (Rel. $L_{2}$ Err). The best model is bold, the second best is red, and the third best is blue.
@@
  • Regarding the Izhikevich model, RNNs (GRU and LSTM) perform the best, while Mamba is the third best.
  • Regarding the LIF model, Mamba performs the second best. However, the best model, Oformer with vanilla attention (Oformer-V), is much more costly than Mamba due to the long sequence with 200 time steps. Hence, Mamba achieves good results in a more reasonable time, which is 4x shorter than Oformer-V. Mamba can capture discontinuous dynamics.
  • Overall, due to the longer sequence length with 401 steps in the Izhikevich model and 200 steps in the LIF model, Transformers are very costly, even with linear cost attentions such as Oformer with Galerkin/Fourier attention and GNOT. In contrast, RNNs and Mamba maintain a lower cost and converge faster.
@@
To conclude this computational result, we visualize some representative Mamba’s predictions in Figure 2. The first row shows the Mamba prediction visualization on the Izhikevich model, where Mamba’s relative $L_{2}$ error is 1.327E-1, while the second row shows that on the LIF model, where Mamba’s relative $L_{2}$ error is 1.905E-2. It is obvious that the first row’s errors are larger than the second row’s errors, which is consistent with our quantitative results. The visualization also suggests that the larger error in the Izhikevich may stem from its more stiff peaks compared with the LIF model. Specifically, the discontinuous solution in the Izhikevich model tends to have six peaks, while the LIF model only has two to three peaks. The visualization results further confirm Mamba’s capability to capture rough solutions in ODE operators with finite regularity.

### v2

@@ 2.1 Neural Operators
Neural operators are a special class of machine learning models designed to learn (nonlinear) mappings between infinite-dimensional spaces, making them particularly suitable for efficiently solving partial differential equations (PDEs). The dynamical systems problem under consideration in this paper also falls into this field. Unlike traditional numerical methods that require mesh generation and can be computationally intensive, neural operators learn a representation of the PDE solution operator directly, allowing for fast evaluations of new inputs once trained. This approach provides significant advantages in terms of generalization and scalability, enabling more rapid and flexible solutions across a variety of complex scenarios in physics and engineering.
@@ 3.2 State Space Model (SSM) and Mamba
where $\boldsymbol{\bar{A}}:=\exp(\Delta\cdot\boldsymbol{A})$, $\boldsymbol{\bar{B}}:=(\Delta\cdot\boldsymbol{A})^{-1}(\exp(\Delta\cdot\boldsymbol{A})-I)\cdot\Delta\boldsymbol{B}$, and $\Delta$ is the discretization step size. The recurrence formulas (2) inherently enable the modeling of unbounded context and linear-time inference, but they are not conducive to efficient parallel training. Fortunately, equation (2) can be further rewritten using a convolution formula as
@@ 3.3 Transformers
To address the efficiency limitation, several attention variants with a linear complexity have been recently proposed and studied. In the context of operator learning, a Galerkin-type attention was first introduced by [5] as a linear variant of attention and later employed by [28, 20],
@@ 4 Computational Experiments
We describe the implementation in detail. We generate 10K train, validation, and test data with a sensor/discretization size of 100, i.e., the sensors are on $\{0.01,0.02,\cdots,0.99,1\}$, where the input function $u(t)$ and the target function $s(t)/s_{t}(t)$ are discretized. Following the original setting in DeepONet, the input functions $u(t)$ are randomly drawn from a mean zero Gaussian random field (GRF) with a length scale of 0.2, i.e., the input functions follow the distribution $\mathcal{GP}\left(0,k_{l}(x,y)\right)$ where $k_{l}(x,y)=\exp\left(-\frac{(x-y)^{2}}{2l^{2}}\right)$ is the RBF kernel with length scale $l$ and $\mathcal{GP}$ denotes a Gaussian process. The output function $s(t)$ or $s_{1}(t)$ can correspondingly be solved via traditional numerical methods. The train loss and the test metric are all the mean square errors (MSE). We train all models for 10,001 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [27] with a 1e-3 initial learning rate, which decays linearly to zero at the end of the training. We keep the parameter numbers of different models similar for a fair comparison. We run the code for five independent random seeds and report the average.
The model structure hyperparameter settings are summarized below. We set an approximate parameter budget of 10,000. The time series model search space is model depth in $\{1,2,3\}$, and hidden dimensions satisfying the budget. We set all Transformer-related models’ head numbers as four and the number of experts in GNOT as two. We set the Transformer and Mamba feedforward dimensions as their respective hidden dimension. We opt for one-layer models with various hidden dimensions to satisfy the budget for all time series models, including LSTM, GRU, Mamba, and Transformers. Specifically, LSTM, GRU, Mamba, and Mamba2 are one-layer models with 32 hidden dimensions. Transformer [52] has 40 hidden and feedforward dimensions, four heads, and one layer. For all three Oformers with vanilla, Galerkin, or Fourier attention [28], we opt for one-layer Oformers with 24 hidden and feedforward dimensions and four heads. GNOT has one layer, 16 hidden and feedforward dimensions, four heads, and two experts. We set the same depth for these time series models for a fair comparison, and we use the validation dataset to verify that shallow and wide time series models are better than deep ($2\sim 3$ layers) and narrow ones in terms of both speed and error. Next, we report the hyperparameter choice on NOs. We keep DeepONet’s trunk and branch nets as MLP with the same structure, whose depth is $2\sim 5$. We find the deepest model is the best, i.e., DeepONet’s [33] trunk and branch nets are all five-layer networks with 24 hidden dimensions. FNO [29] is a four-layer model with 16 hidden dimensions and eight modes. We validate other FNO models with $2\sim 4$ layers, $8/16$ modes, and a corresponding hidden dimension to satisfy the parameter count budget and choose the best one. We follow Cao et al. [4] and choose a one-layer LNO following their experiments on 1D dynamical systems. LNO has 16 hidden dimensions and eight modes. We validate other LNO models with one layer, $4/8/16$ modes, and a corresponding hidden dimension to satisfy the parameter count budget and choose the best one.
@@
LSTM 8545 1643MiB 31min 1.625E-7$\pm$1.185E-7 9.509E-6$\pm$8.892E-6 1.322E-8$\pm$9.719E-9
GRU 6433 1629MiB 30min 3.627E-8$\pm$2.771E-8 3.215E-6$\pm$2.276E-6 7.117E-9$\pm$1.953E-9
DeepONet 7272 1381MiB 34min 1.296E-7$\pm$4.094E-8 2.147E-5$\pm$1.435E-6 3.500E-8$\pm$3.479E-8
FNO 9329 1833MiB 73min 1.536E-4$\pm$6.927E-6 6.519E-4$\pm$1.032E-4 4.008E-5$\pm$8.730E-6
LNO 6721 1797MiB 52min 2.214E-7$\pm$1.928E-7 4.775E-5$\pm$1.944E-5 1.073E-8$\pm$4.329E-9
Transformer 10201 1443MiB 49min 5.197E-5$\pm$2.183E-5 4.314E-4$\pm$1.098E-5 6.786E-6$\pm$7.092E-7
Oformer-V 11209 1439MiB 41min 4.210E-8$\pm$1.139E-8 1.352E-6$\pm$6.782E-7 7.549E-9$\pm$4.378E-9
Oformer-G 11161 1379MiB 49min 3.472E-8$\pm$3.089E-9 6.036E-7$\pm$1.256E-7 4.582E-9$\pm$1.115E-9
Oformer-F 11161 1379MiB 49min 3.598E-8$\pm$6.595E-9 4.582E-7$\pm$9.735E-8 4.302E-9$\pm$1.538E-9
GNOT 11187 1403MiB 91min 4.155E-8$\pm$7.239E-9 6.253E-7$\pm$2.714E-7 5.547E-9$\pm$1.246E-9
Mamba 6593 1425MiB 35min 3.333E-9$\pm$1.288E-9 4.351E-8$\pm$1.782E-8 1.301E-9$\pm$3.273E-10
Mamba2 6009 1282MiB 109min 2.106E-8$\pm$8.198E-9 2.015E-7$\pm$9.302E-8 4.464E-8$\pm$3.981E-8
Table 2: Test MSE error results for 1D dynamical systems, including the anti-derivative operator, nonlinear ODE, and gravity pendulum, from DeepONet’s benchmark cases [33] corresponding to Section 4.1 in this paper. The errors are MSE following Lu et al. [33]. Optimization trajectories are visualized in Figure 2. The best model is bold, the second best is red, and the third best is blue.
Figure 2: Visualization of various models’ loss trajectories for the 1D Pendulum system following Lu et al. [33] corresponding to Section 4.1 in this paper. The detailed quantitative results for this setting are presented in Table 2. Subfigure (a) is Mamba versus RNNs; subfigure (b) is Mamba versus NOs; and subfigure (c) is Mamba versus Transformers.
Table 2 shows the computational results, where we report the model name, the number of total parameters in the model, the memory cost in Mebibyte (MiB) obtained from the command “nvidia-smi” on the Linux GPU server where only the training part is considered into the memory cost evaluation, the total training time in minutes (min), and the model MSE prediction error on the three operator learning tasks. The MSE error metric was initially adopted by Lu et al. [33], and we strictly follow them for a clear comparison.
@@
  • Regarding the proposed Mamba’s performances, Mamba is the best model for all three operator learning tasks. It outperforms the second-best model by a large margin. Mamba2 is generally good and ranks the second best in the anti-derivative and nonlinear ODE operators. Even with a limited 10,000 parameter count budget, Mamba can already achieve a low error.
@@
  • We construct models with a similar number of parameters. Hence, their memory costs are similar. However, the total training time differs drastically due to different models’ computational complexity. Specifically, the expensive self-attention mechanism makes Transformer-related models (Transformer, Oformer, and GNOT) the slowest among time series models, while RNNs and Mamba are relatively faster.
  • The training loss trajectory in the Pendulum problem is visualized in Figure 2. Mamba always maintains its advantage in faster convergence throughout the optimization. In later studies in Sections 4.5 and 4.7, we adopt the same ODE data and Mamba model and report some relative error results, which is at the scale of $10^{-4}\sim 10^{-3}$ for in-distribution IID results. This demonstrates Mamba’s outstanding representation ability under a limited parameter count budget, which does not require larger model sizes for low errors.
@@ 4.2 Finite Regularity and Discontinuous Solutions
The training loss is MSE, and the test metric is the relative $L_{2}$ error. The test metric is similar to the one used by Shih et al. [46] for a direct comparison. We train all models for 10,001 epochs. In each epoch, we traverse the dataset with a minibatch size of 128. We use the Adam optimizer [27] with a 1e-3 initial learning rate, which decays linearly to zero at the end of training. We keep the number of parameters of different models similar for a fair comparison. We run the code for five independent random seeds and report the average.
@@
The model structure hyperparameter settings are summarized below. During validation, we train the model on 80% of the training set and validate on the rest 20%. The model search space is the same as the previous Section 4.1, and we find the following outperforming model structures. LSTM and GRU are one-layer models with 32 hidden dimensions. DeepONet’s trunk and branch nets are all five-layer networks with 24 hidden dimensions. FNO is a four-layer model with 16 hidden dimensions and eight modes. We choose a one-layer LNO with eight hidden dimensions and four modes. We opt for a one-layer Transformer with 40 hidden and feedforward dimensions and four heads. For all three Oformers with vanilla, Galerkin, or Fourier attention, we opt for one-layer Oformers with 24 hidden and feedforward dimensions and four heads. GNOT has three layers, 16 hidden and feedforward dimensions, four heads, and two experts. Regarding Mamba, we opt for a block with skip connection and layer normalization, with 16 hidden and intermediate dimensions. Again, for all time series models, including RNNs, Mamba, and Transformers, we keep the same one-layer structure and similar parameter count budgets for a fair comparison.
The Izhikevich Model Tempered Fractional LIF Model
Model Params Memory Time Relative $L_{2}$ Error Model Params Memory Time Relative $L_{2}$ Error
LSTM 8545 1863MiB 10min 1.225E-1$\pm$9.256E-3 LSTM 8577 1707MiB 7min 2.181E-2$\pm$9.073E-5
GRU 6433 1821MiB 10min 1.138E-1$\pm$8.422E-3 GRU 6465 1703MiB 7min 2.056E-2$\pm$1.556E-4
DeepONet 14496 1383MiB 9min 2.150E-1$\pm$1.155E-4 DeepONet 14472 1383MiB 7min 2.782E-1$\pm$3.581E-3
FNO 9329 1859MiB 17min 1.517E-1$\pm$2.508E-4 FNO 9345 1839MiB 15min 2.203E-2$\pm$3.385E-4
LNO 1889 1917MiB 12min 1.786E-1$\pm$1.842E-2 LNO 1897 1849MiB 9min 2.635E-1$\pm$3.198E-2
Mamba 9649 1599MiB 12min 1.359E-1$\pm$1.421E-2 Mamba 9665 1493MiB 8min 2.132E-2$\pm$3.087E-4
Transformer 10201 2951MiB 44min 2.274E-1$\pm$1.012E-2 Transformer 10241 1735MiB 13min 4.549E-2$\pm$2.509E-3
Oformer-V 11209 2743MiB 39min 2.325E-1$\pm$2.189E-2 Oformer-V 11233 1711MiB 12min 2.395E-2$\pm$5.660E-4
Oformer-G 11161 1579MiB 31min 2.243E-1$\pm$2.541E-2 Oformer-G 11185 1443MiB 14min 6.419E-2$\pm$9.055E-3
Oformer-F 11161 1579MiB 31min 2.202E-1$\pm$2.491E-2 Oformer-F 11185 1443MiB 14min 6.623E-2$\pm$7.160E-3
GNOT 11187 1631MiB 61min 1.782E-1$\pm$2.864E-2 GNOT 11203 1445MiB 27min 6.262E-2$\pm$1.207E-2
@@
Table 3: Results for finite regularity and discontinuous solutions adopted from Shih et al. [46] corresponding to Section 4.2 in this paper. Visualization of the solutions is presented in Figure 3. Here, we present the test relative $L_{2}$ errors. The best model is bold, the second best is red, and the third best is blue. Some loss trajectories are visualized in Figure 4.
4.2.1 Main Results
@@
  • Regarding the Izhikevich model, RNNs (GRU and LSTM) perform the best, while Mamba is the third best. Regarding the LIF model, Mamba performs the second best. Hence, Mamba can capture discontinuous dynamics.
  • Overall, due to the longer sequence length with 401 steps in the Izhikevich model and 200 steps in the LIF model, Transformers are more costly, even with linear cost attentions such as Oformer with Galerkin/Fourier attention and GNOT. In contrast, RNNs and Mamba maintain a lower cost and converge faster.
@@
We visualize some representative Mamba’s predictions in Figure 3. The first row shows the Mamba prediction visualization on the Izhikevich model, where Mamba’s relative $L_{2}$ error is 1.257E-1, while the second row shows that on the LIF model, where Mamba’s relative $L_{2}$ error is 2.126E-2. It is obvious that the first row’s errors are larger than the second row’s errors, which is consistent with our quantitative results. The visualization also suggests that the larger error in the Izhikevich may stem from its more stiff peaks compared with the LIF model. Specifically, the discontinuous solution in the Izhikevich model tends to have six peaks, while the LIF model only has two to three peaks. The visualization results further confirm Mamba’s capability to capture rough solutions in ODE operators with finite regularity.
Figure 4: Visualization of top three models’ loss trajectories for operators with finite regularity following Shih et al. [46] corresponding to Section 4.2 in this paper. The detailed quantitative results for this setting are presented in Table 3. Subfigure (a) is the Izhikevich model; subfigure (b) is the LIF model.
To further understand and compare the top three models, namely GRU, LSTM, and Mamba, we plot their respective optimization trajectories on the two problems in Figure 4. Since the LIF model has two to three peaks and is easier to learn in terms of model accuracy, the three models’ train loss decreases continuously throughout the optimization in (b). However, due to the stronger discontinuity in the Izhikevich model in (a), LSTM and GRU exhibit unstable optimization with spikes in loss during training, while Mamba remains stable. Despite Mamba’s slightly larger error, it demonstrates better training stability under solution discontinuity in operator learning.
4.2.2 Additional Study: Larger Model for Better Performances
Due to the discontinuous solutions, the relative $L_{2}$ test error results for the Izhikevich model are larger than $10^{-1}$, which is not satisfactory. Hence, we will conduct an additional study to achieve better performance with larger model sizes and further understand different models’ accuracy under larger parameter counts. We selected the top three models for this study: LSTM, GRU, and Mamba, with various depths and hidden dimensions. We keep other settings the same except for controlling the model size. The additional results are shown in Table 4. The lowest error is achieved by Mamba with three layers and 16 hidden dimensions, demonstrating Mamba’s advantage under various model sizes.
Additional Study on the Izhikevich Model
Model Depth Width Params Memory Time Relative $L_{2}$ Error
LSTM 1 32 8545 1863MiB 10min 1.225E-1$\pm$9.256E-3
2 32 16993 1885MiB 13min 1.240E-1$\pm$2.110E-2
3 32 25441 1923MiB 15min 1.218E-1$\pm$6.069E-3
1 64 33473 2175MiB 11min 1.216E-1$\pm$1.190E-2
GRU 1 32 6433 1821MiB 10min 1.138E-1$\pm$8.422E-3
2 32 12769 1861MiB 14min 1.111E-1$\pm$5.581E-3
3 32 19105 1905MiB 16min 1.193E-1$\pm$1.792E-2
1 64 25153 2087MiB 12min 1.249E-1$\pm$1.497E-2
Mamba 1 16 9649 1599MiB 12min 1.359E-1$\pm$1.421E-2
2 16 19217 1759MiB 18min 1.175E-1$\pm$6.209E-3
3 16 28785 1919MiB 28min 1.045E-1$\pm$2.698E-3
1 32 22497 1655MiB 14min 1.305E-1$\pm$5.578E-3
Table 4: Results for larger models on the Izhikevich model corresponding to Section 4.2.2.